# Reinforcement Fine-Tuning Naturally Mitigates Forgetting in Continual Post-Training

### v3

@@ Abstract
Continual post-training (CPT) is a popular and effective technique for adapting foundation models like multimodal large language models to specific and ever-evolving downstream tasks. While existing research has primarily concentrated on methods like data replay, model expansion, or parameter regularization, the fundamental role of the learning paradigm within CPT remains largely unexplored. This paper presents a comparative analysis of two core post-training paradigms: supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT), investigating their respective impacts on knowledge retention during CPT. Our experiments are conducted on a benchmark comprising seven diverse multimodal tasks, utilizing Qwen2.5-VL-7B-Instruct as the base model for continual post-training. The investigation yields two significant findings: (1) When continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks. In contrast, RFT inherently preserves prior knowledge and achieve performance comparable to multi-task training. (2) RFT successfully protects and even enhances the model’s general knowledge on standard benchmarks (e.g., MMMU and MMLU-Pro). Conversely, SFT degrades general model capabilities severely. Further analysis reveals that this stability is not primarily due to explicit mechanisms like KL penalty or chain-of-thought reasoning. Instead, we identify an implicit regularization mechanism inherent to RFT as a key contributing factor. Our theoretical analysis suggests that RFT’s gradient updates are naturally scaled by the reward variance, acting as a data-dependent regularizer that inherently protects previously acquired knowledge. Finally, we propose a rollout-based instance filtering algorithm to enhance the stability and efficiency of RFT. Our comprehensive study demonstrates the superiority of RFT as a robust paradigm for continual post-training.  [^1] Our code is provided in the supplementary material. An anonymous link for review is: https://anonymous.4open.science/r/RFTvsSFT-A999
@@ 1 Introduction
Recent advancements in multimodal large language models (MLLMs) have demonstrated remarkable capabilities in complex world understanding (Achiam et al., 2023; Liu et al., 2024; Wang et al., 2024a). To align with the demands of real-world deployment, MLLMs must adapt to a stream of data and evolving user requirements, incorporating new skills and domain knowledge over time (Zhu et al., 2024). This calls for an efficient and scalable continual post-training (CPT) paradigm. A key challenge in CPT is the well-known phenomenon of catastrophic forgetting (McCloskey & Cohen, 1989), where adapting to a new task leads to a severe degradation of performance on previously learned tasks. To reduce forgetting, recent studies (Guo et al., 2025c) focus on data replay (Maharana et al., 2025; Lee et al., 2025; Wang et al., 2025), model expansion (Zhao et al., 2025; Guo et al., 2025b; Zeng et al., 2024), and explicit regularization (Liu et al., 2025a). Nevertheless, existing methods typically leverage the supervised fine-tuning (SFT) paradigm by default, and the role of the fundamental fine-tuning paradigm in CPT has been overlooked.
Recently, reinforcement fine-tuning (RFT), which optimizes models based on feedback from generated outputs, has significantly advanced foundation model post-training (Chu et al., 2025; Shao et al., 2024; Guo et al., 2025a). To the best of our knowledge, this work presents the first direct comparative investigation into whether SFT or RFT is the more suitable paradigm for CPT, focusing on knowledge preservation for both specific downstream tasks and general capabilities. Experimentally, we continually fine-tune the Qwen2.5-VL-7B-Instruct model (Bai et al., 2025b) on a benchmark comprising diverse multimodal tasks covering various domains. To fully reflect the knowledge preservation ability, we evaluate forgetting on both learned specific tasks and general benchmarks such as MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a).
@@
To understand how RFT mitigates forgetting during CPT, we conduct additional experiments with the popular and representative group relative policy optimization (GRPO) framework (Shao et al., 2024). We analyze the impact of KL divergence penalty and chain-of-thought (CoT) reasoning (Wei et al., 2022) on forgetting mitigation. Particularly, the KL divergence penalty prevents the policy from changing too drastically, similar to the well-known knowledge distillation in continual learning (Li & Hoiem, 2017). However, our analysis indicates that these explicit mechanisms are not the primary drivers of forgetting mitigation. We instead attribute this phenomenon to an implicit regularization effect within RFT. We offer a theoretical perspective suggesting that RFT’s updates are inherently more conservative in parameter subspaces sensitive to prior tasks. This conservatism is naturally scaled by the variance of the reward signal, creating a data-dependent regularization that dampens updates on uncertain samples, thus protecting established knowledge. Last but not least, we observe that the learning process of RFT can be highly inefficient. Thus, we introduce a rollout-based instance filtering algorithm that enhances the stability of GRPO while still being an excellent knowledge protector.
@@ 2 Related Works
Post-training is a critical stage for refining the capabilities of pre-trained foundation models (Shao et al., 2024; Chu et al., 2025; Achiam et al., 2023). SFT on task-specific or instruction-formatted datasets is a common approach to adapt models to downstream applications (Chung et al., 2024; Zhou et al., 2023). For example, Chung et al. (2024) demonstrated that by scaling the number of tasks and model size, and incorporating CoT data, SFT significantly enhances the performance and generalization of various large language models across diverse benchmarks. Recently, RFT has gained prominence for aligning models with human preferences or improving performance on specific objectives (Liu et al., 2025c; Zhai et al., 2024; Shao et al., 2024; Luong et al., 2024; Li et al., 2025; 2023c; Ahmadian et al., 2024). Particularly, GRPO (Shao et al., 2024) largely enhances mathematical reasoning and optimizes memory usage, being a popular method for post-training of large language models. Liu et al. (2025b) revealed inherent biases in the GRPO algorithm, then introduces an unbiased optimization method that improves token efficiency while maintaining reasoning performance. Visual-RFT (Liu et al., 2025c) boosts MLLMs by using reinforcement learning with rule-based visual rewards, making them more data-efficient and better at various visual tasks than traditional SFT. Recently, Chu et al. (2025) demonstrated that reinforcement learning significantly enhances the generalization capabilities of foundation models, while SFT primarily leads to memorization. In this work, we study the comparative effect of SFT and RFT on knowledge retention in MLLMs continual post-training.
@@ 3 Preliminaries
where $\beta>0$. ReMax (Li et al., 2023c) use the reward of a greedy decoding response $a$ as a baseline. For a single sampled response $a$, the objective is to maximize:
$\mathcal{J}_{\text{ReMax}}(\theta)=\mathbb{E}_{x,a\sim\pi_{\theta}}\left[(r(x,a)-r(x,a))\log\pi_{\theta}(a|x)\right].$ (4)
@@ 4.1 Experimental Setup
We adopt the open-source Qwen2.5-VL-7B-Instruct (Bai et al., 2025a) as our base model, primarily due to its demonstrated superiority in vision-language comprehension and its favorable resource footprint, which is crucial for practical deployment. We continually fine-tune the model on diverse vision-language datasets (ScienceQA (Saikh et al., 2022), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), GQA (Hudson & Manning, 2019), Geometry3K (Lu et al., 2021), PathVQA (He et al., 2020), Super-CLEVR (Li et al., 2023b)), covering a wide range of common downstream applications. After the end of CPT, evaluation is performed on the test sets of all previously encountered tasks. Additionally, to fully assess the knowledge preservation ability, we evaluate the model on diverse, general benchmarks at the end of learning all downstream tasks. Specifically, we evaluate the model on three specialized benchmarks: MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a). Particularly, we include POPE to systematically assess whether CPT induces object hallucination in MLLMs. A detailed description of those datasets is provided in the Appendix.
@@
All experiments employ full-parameter fine-tuning for both SFT and RFT to ensure comprehensive capability assessment. Experiments of SFT are conducted using the llamafactory (Zheng et al., 2024) framework, with a learning rate of 1e51e-5 and a batch size of 24. RFT methods (GRPO, ReMax, and RLOO) are implemented using the easyR1 (Zheng et al., 2025) framework, building upon Verl (Sheng et al., 2024). A consistent configuration is applied across RFT methods to ensure an equitable comparison: a learning rate of 1e61e-6, a rollout batch size of 512, a sampling temperature of 1.0, with KL-divergence coefficient $\beta=0.01$. Specifically, GRPO is implemented adhering to its foundational methodology, with a group size set to 8. ReMax followed its core algorithm, and RLOO adopted the official Hugging Face algorithm.
@@ 5.1 The Roles of CoT and KL Penalty
We test two primary hypotheses: (1) The KL-divergence penalty, by regularizing policy updates, acts as a form of knowledge distillation (Li & Hoiem, 2017) that preserves past knowledge. (2) The complex reasoning structure of CoT builds more abstract and resilient knowledge representations, protecting them from being overwritten. Thus, we evaluate three GRPO variants against the SFT baseline: GRPO w/o KL: trained with CoT prompts but without the KL penalty term. GRPO w/o CoT: trained without CoT prompts, using direct question-answering format but retaining the KL penalty.
@@
Figure 4: Loss landscapes illustration of SFT and RFT based continual post-training.
@@ 5.2 Implicit Regularization from Reward Variance
The full proof is provided in Appendix B. Proposition 5.2 offers an intuition: the expected impact of an RFT update on prior knowledge is not fixed but is dynamically scaled by the reward variance. For an uncertain sample where the model generates diverse responses with high reward variance, the update magnitude in sensitive directions is naturally dampened, thus protecting established knowledge. Conversely, for samples where the model produces consistently high-reward responses (low variance), the update is more aggressive. This inherent, data-dependent regularization mechanism contrasts sharply with SFT’s uniform, high-variance gradients, offering a compelling explanation for the stability observed in our experiments and illustrated in Figure 4.
@@ 5.3 RIF-RFT: Enhancing Stability and Efficiency of RFT
Our analysis in Section 5.2 reveals that RFT’s resilience to forgetting is based in a reward-variance-scaled regularization. However, this mechanism’s effectiveness relies on the model’s ability to generate responses that produce a meaningful reward signal. We identify a critical failure mode when the model is faced with incompetent samples: training instances for which the current policy $\pi_{\theta}$ consistently fails to produce non-zero rewarded outputs. For such samples, the advantage estimates $A(x,a)$ are dominated by noise, leading to high-variance gradient estimates.
To address this challenge, we propose a simple yet effective method: Rollout-based Instance Filtering for RFT (RIF-RFT). The motivation, illustrated in Figure 4, is to prune the training data by identifying and discarding these incompetent samples before the RFT training. By filtering them out, RFT focuses its capacity on instances where it can receive a productive learning signal, stabilizing the regularization effect and improving efficiency.
The mechanism is formalized in Algorithm 1 in Appendix C. For each instance in a new task’s dataset $\mathcal{D}_{k}$, we perform a small number of policy rollouts. If at least one of these rollouts produces a response with a reward greater than a minimal threshold $\tau$, we classify the instance and retain it in $\mathcal{D}_{k}^{\text{filt}}$. As shown in Table 5, RIF-RFT achieves competitive performance (57.5% of AvgAcc) and strong knowledge retention (-4.5% of FM), which is superior to SFT with a FM of -10.4%. Meanwhile, RIF-RFT use only 48.4% of training data. This demonstrates a compelling trade-off between efficiency and robustness.
@@ Appendix A Dataset Information
  • GQA (Hudson & Manning, 2019) is designed for compositional question answering over real-world images with a strong emphasis on spatial understanding and object relationships.
@@ Appendix C Pseudo Code of RIF-RFT
   Initialize $R_{\max}\leftarrow$-$\infty$;
@@
      Update: $R_{\max}\leftarrow\max(R_{\max},R(y_{ij}))$;
 if $R_{\max}>\tau$ then
@@
Appendix D Statement on the Use of LLM
We disclose that LLMs were used to assist with writing. Their role was limited to improving the grammar and overall readability of the manuscript. The core research ideas, experimental results, and scientific claims presented are the work of the authors.

### v4

@@ Abstract
Continual post-training (CPT) is a popular and effective technique for adapting foundation models like multimodal large language models to specific and ever-evolving downstream tasks. While existing research has primarily concentrated on methods like data replay, model expansion, or parameter regularization, the fundamental role of the learning paradigm within CPT remains largely unexplored. This paper presents a comparative analysis of two core post-training paradigms: supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT), investigating their respective impacts on knowledge retention during CPT. Our experiments are conducted on a benchmark comprising seven diverse multimodal tasks, utilizing Qwen2.5-VL-7B-Instruct as the base model for continual post-training. The investigation yields two significant findings: (1) When continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks. In contrast, RFT inherently preserves prior knowledge and achieve performance comparable to multi-task training. (2) RFT successfully protects and even enhances the model’s general knowledge on standard benchmarks (e.g., MMMU and MMLU-Pro). Conversely, SFT degrades general model capabilities severely. Further analysis reveals that this stability is not primarily due to explicit mechanisms like KL penalty or chain-of-thought reasoning. Instead, we identify an implicit regularization mechanism inherent to RFT as a key contributing factor. Our theoretical analysis suggests that RFT’s gradient updates are naturally scaled by the reward variance, acting as a data-dependent regularizer that inherently protects previously acquired knowledge. Finally, we propose a rollout-based instance filtering algorithm to enhance the stability and efficiency of RFT. Our comprehensive study demonstrates the superiority of RFT as a robust paradigm for continual post-training.  [^1] Our code is provided in the supplementary material. An anonymous link for review is: https://github.com/zhhvvv/rft_vs_sft
@@ 1 Introduction
Recent advancements in multimodal large language models (MLLMs) have demonstrated remarkable capabilities in complex world understanding (Achiam et al., 2023; Liu et al., 2024; Wang et al., 2024a). To align with the demands of real-world deployment, MLLMs must adapt to a stream of data and evolving user requirements, incorporating new skills and domain knowledge over time (Zhu et al., 2024). This calls for an efficient and scalable continual post-training (CPT) paradigm. A key challenge in CPT is the well-known phenomenon of catastrophic forgetting (McCloskey and Cohen, 1989), where adapting to a new task leads to a severe degradation of performance on previously learned tasks. To reduce forgetting, recent studies (Guo et al., 2025c) focus on data replay (Maharana et al., 2025; Lee et al., 2025; Wang et al., 2025), model expansion (Zhao et al., 2025; Guo et al., 2025b; Zeng et al., 2024), and explicit regularization (Liu et al., 2025a). Nevertheless, existing methods typically leverage the supervised fine-tuning (SFT) paradigm by default, and the role of the fundamental fine-tuning paradigm in CPT has been overlooked.
Recently, reinforcement fine-tuning (RFT), which optimizes models based on feedback from generated outputs, has significantly advanced foundation model post-training (Chu et al., 2025; Shao et al., 2024; Guo et al., 2025a). To the best of our knowledge, this work presents the first direct comparative investigation into whether SFT or RFT is the more suitable paradigm for CPT, focusing on knowledge preservation for both specific downstream tasks and general capabilities. Experimentally, we continually fine-tune the Qwen2.5-VL-7B-Instruct model (Bai et al., 2025a) on a benchmark comprising diverse multimodal tasks covering various domains. To fully reflect the knowledge preservation ability, we evaluate forgetting on both learned specific tasks and general benchmarks such as MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a).
@@
To understand how RFT mitigates forgetting during CPT, we conduct additional experiments with the popular and representative group relative policy optimization (GRPO) framework (Shao et al., 2024). We analyze the impact of KL divergence penalty and chain-of-thought (CoT) reasoning (Wei et al., 2022) on forgetting mitigation. Particularly, the KL divergence penalty prevents the policy from changing too drastically, similar to the well-known knowledge distillation in continual learning (Li and Hoiem, 2017). However, our analysis indicates that these explicit mechanisms are not the primary drivers of forgetting mitigation. We instead attribute this phenomenon to an implicit regularization effect within RFT. We offer a theoretical perspective suggesting that RFT’s updates are inherently more conservative in parameter subspaces sensitive to prior tasks. This conservatism is naturally scaled by the variance of the reward signal, creating a data-dependent regularization that dampens updates on uncertain samples, thus protecting established knowledge. Last but not least, we observe that the learning process of RFT can be highly inefficient. Thus, we introduce a rollout-based instance filtering algorithm that enhances the stability of GRPO while still being an excellent knowledge protector.
@@ 2 Related Works
Post-training is a critical stage for refining the capabilities of pre-trained foundation models (Shao et al., 2024; Chu et al., 2025; Achiam et al., 2023). SFT on task-specific or instruction-formatted datasets is a common approach to adapt models to downstream applications (Chung et al., 2024; Zhou et al., 2023). For example, Chung et al. (2024) demonstrated that by scaling the number of tasks and model size, and incorporating CoT data, SFT significantly enhances the performance and generalization of various large language models across diverse benchmarks. Recently, RFT has gained prominence for aligning models with human preferences or improving performance on specific objectives (Liu et al., 2025c; Zhai et al., 2024; Shao et al., 2024; Luong et al., 2024; Li et al., 2025; 2023c; Ahmadian et al., 2024). Particularly, GRPO (Shao et al., 2024) largely enhances mathematical reasoning and optimizes memory usage, being a popular method for post-training of large language models. Liu et al. (2025b) revealed inherent biases in the GRPO algorithm, then introduces an unbiased optimization method that improves token efficiency while maintaining reasoning performance. Visual-RFT (Liu et al., 2025c) boosts MLLMs by using reinforcement learning with rule-based visual rewards, making them more data-efficient and better at various visual tasks than traditional SFT. Recently, Chu et al. (2025) demonstrated that reinforcement learning significantly enhances the generalization capabilities of foundation models, while SFT primarily leads to memorization. In this work, we study the comparative effect of SFT and RFT on knowledge retention in MLLMs continual post-training. Recent work by Zhang et al. (2025) investigates SFT and RFT from a data perspective, showing that incorporating reasoning trajectories in SFT can reduce forgetting. Their findings complement our work by highlighting how data format affects SFT’s stability, while we demonstrate that RFT provides inherent forgetting mitigation without reasoning format. Together, these studies provide comprehensive guidance for post-training paradigm selection.
@@ 3 Preliminaries
where $\beta>0$. ReMax (Li et al., 2023c) use the reward of a greedy decoding response $\hat{a}$ as a baseline. For a single sampled response $a$, the objective is to maximize:
$\mathcal{J}_{\text{ReMax}}(\theta)=\mathbb{E}_{x,a\sim\pi_{\theta}}\left[(r(x,a)-r(x,\hat{a}))\log\pi_{\theta}(a|x)\right].$ (4)
@@ 4.1 Experimental Setup
We adopt the open-source Qwen2.5-VL-7B-Instruct (Bai et al., 2025b) as our base model, primarily due to its demonstrated superiority in vision-language comprehension and its favorable resource footprint, which is crucial for practical deployment. We continually fine-tune the model on diverse vision-language datasets (ScienceQA (Saikh et al., 2022), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), GQA (Hudson and Manning, 2019), Geometry3K (Lu et al., 2021), PathVQA (He et al., 2020), Super-CLEVR (Li et al., 2023b)), covering a wide range of common downstream applications. After the end of CPT, evaluation is performed on the test sets of all previously encountered tasks. Additionally, to fully assess the knowledge preservation ability, we evaluate the model on diverse, general benchmarks at the end of learning all downstream tasks. Specifically, we evaluate the model on three specialized benchmarks: MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a). Particularly, we include POPE to systematically assess whether CPT induces object hallucination in MLLMs. A detailed description of those datasets is provided in the Appendix A.
@@
All experiments employ full-parameter fine-tuning for both SFT and RFT to ensure comprehensive capability assessment. Experiments of SFT are conducted using the llamafactory (Zheng et al., 2024) framework, with a learning rate of 1e51e-5 and a batch size of 24. RFT methods (GRPO, ReMax, and RLOO) are implemented using the easyR1 (Zheng et al., 2025) framework, building upon Verl (Sheng et al., 2024). A consistent configuration is applied across RFT methods to ensure an equitable comparison: a learning rate of 1e61e-6, a rollout batch size of 512, a sampling temperature of 1.0, with KL-divergence coefficient $\beta=0.01$. Specifically, GRPO is implemented adhering to its foundational methodology, with a group size set to 8. ReMax followed its core algorithm, and RLOO adopted the official Hugging Face algorithm. To ensure the generality of our findings, we conduct additional experiments across different model architectures, scales, and task domains, with detailed results provided in Appendix D.
@@ 5.1 The Roles of CoT and KL Penalty
We test two primary hypotheses: (1) The KL-divergence penalty, by regularizing policy updates, acts as a form of knowledge distillation (Li and Hoiem, 2017) that preserves past knowledge. (2) The complex reasoning structure of CoT builds more abstract and resilient knowledge representations, protecting them from being overwritten. Thus, we evaluate three GRPO variants against the SFT baseline: GRPO w/o KL: trained with CoT prompts but without the KL penalty term. GRPO w/o CoT: trained without CoT prompts, using direct question-answering format but retaining the KL penalty.
@@
Figure 4: Only samples producing non-zero reward variance yield effective policy gradients; RIF-RFT improves sample efficiency by focusing training on such informative samples.
@@ 5.2 Implicit Regularization from Reward Variance
The full proof is provided in Appendix B. Proposition 5.2 offers an intuition: the expected impact of an RFT update on prior knowledge is not fixed but is dynamically scaled by the reward variance. For an uncertain sample where the model generates diverse responses with high reward variance, the update magnitude in sensitive directions is naturally dampened, thus protecting established knowledge. Conversely, for samples where the model produces consistently high-reward responses, the update is more aggressive. This inherent, data-dependent regularization mechanism contrasts with SFT’s uniform, high-variance gradients, offering a compelling explanation for the stability observed in our experiments and illustrated in Figure 4.
@@ 5.3 RIF-RFT: Enhancing Stability and Efficiency of RFT
Our analysis in Section 5.2 reveals that RFT’s resilience to forgetting is based in a reward-variance-scaled regularization. However, this mechanism’s effectiveness relies on the model’s ability to generate responses that produce a meaningful reward signal. We identify a critical failure mode when the model is faced with incompetent samples: training instances for which the current policy $\pi_{\theta}$ consistently fails to produce non-zero rewarded outputs. For such samples, the advantage estimates $A(x,a)$ collapse to zero or are dominated by noise, yielding no effective policy gradient. This reduces sample efficiency without contributing to meaningful learning.
To address this challenge, we propose a simple yet effective method: Rollout-based Instance Filtering for RFT (RIF-RFT). The motivation is to prune the training data by identifying and discarding these incompetent samples before the RFT training. By filtering them out, RFT focuses its capacity on instances where it can receive a productive learning signal, stabilizing the regularization effect and improving efficiency. Note that as training progresses, samples that initially yielded zero reward may become learnable. RIF-RFT trades this marginal adaptability for computational savings.
The mechanism is formalized in Algorithm 1 in Appendix C. For each instance in a new task’s dataset $\mathcal{D}_{k}$, we perform a small number of policy rollouts. If at least one of these rollouts produces a response with a reward greater than a minimal threshold $\tau$, we classify the instance and retain it in $\mathcal{D}_{k}^{\text{filt}}$. As shown in Table 5, while full-data GRPO achieves the best performance, it processes many samples with zero reward variance that yield no effective policy gradients. RIF-RFT addresses this inefficiency by filtering such samples a priori, maintaining strong anti-forgetting properties. This demonstrates a compelling trade-off between efficiency and robustness.
@@ Appendix A Dataset Information
  • GQA (Hudson and Manning, 2019) is designed for compositional question answering over real-world images with a strong emphasis on spatial understanding and object relationships.
@@ Appendix C Pseudo Code of RIF-RFT
   Initialize $R_{\text{sum}}\leftarrow 0$;
@@
      Update: $R_{\text{sum}}\leftarrow R_{\text{sum}}+R(y_{ij})$;
 if $R_{\text{sum}}/N>\tau$ then
@@
Appendix D Robustness and Efficiency Analysis
To validate the generality of our findings beyond the primary Qwen2.5-VL-7B-Instruct model, we conduct extensive experiments across different architectures, model scales, and task domains. These additional studies ensure that the observed forgetting mitigation is an intrinsic property of the RFT paradigm rather than an artifact of a specific model configuration.
D.1 Generalization Across Architectures and Scales
D.1.1 Text-Only Tasks
We first evaluate whether RFT’s forgetting mitigation extends to text-only domains. We utilize the text-only Qwen2.5-7B-Instruct (Yang et al., 2024) and evaluate it on two diverse benchmarks: GSM8K (Cobbe et al., 2021) for mathematical reasoning and USMLE (Jin et al., 2020) for medical knowledge. These tasks provide clear correctness signals suitable for both SFT and RFT paradigms. As shown in Table 6, RFT consistently outperforms SFT. For instance, in the GSM8K $\rightarrow$ USMLE sequence, RFT maintains a forgetting measure (FM) of -1.8%, whereas SFT suffers a significant drop with an FM of -10.4%.
Table 6: Performance evaluation on text-only tasks using Qwen2.5-7B-Instruct.
Method Task Order GSM8K USMLE AvgAcc FM
[…]
D.1.2 Model Scale Analysis
We further evaluate the impact of model scale on forgetting mitigation using Qwen2.5-VL-3B-Instruct (Bai et al., 2025c) and the larger Qwen3-VL-8B-Instruct (Yang et al., 2025) on a subset of our benchmark tasks (sCLEVR, ScienceQA, and TextVQA). The results are summarized in Table 7. We observe that RFT maintains near-zero forgetting across both scales . Notably, the larger 8B model exhibits stronger resilience to catastrophic forgetting under the RFT paradigm compared to the 3B model.
Table 7: Performance comparison across different model scales.
Model Size Method sCLEVR SciQA TextVQA AvgAcc FM
3B GRPO 57.8 92.7 72.8 74.4 -0.4
SFT 51.5 92.3 67.6 70.5 -4.4
8B GRPO 57.0 96.3 76.1 76.5 -0.2
SFT 48.2 91.5 68.9 69.5 -7.1
D.2 Comparison with CL Methods
To compare RFT’s performance against established CL techniques, we compare it with Experience Replay (ER) (Schaul et al., 2015), widely considered one of the most effective baselines. We implement ER with a 25% replay ratio, which represents the upper range suggested by recent work on LLM continual post-training (Abbes et al., 2025). As detailed in Table 8, while ER improves upon vanilla SFT (FM improves from -4.4% to -2.8%), it still lags behind RFT (-0.4%). Furthermore, ER introduces significant storage overhead and potential negative transfer, whereas RFT achieves superior stability inherently without requiring external memory buffers.
Table 8: Comparison between RFT, SFT, and SFT with ER on Qwen2.5-VL-3B-Instruct.
Method sCLEVR SciQA TextVQA AvgAcc FM
SFT 51.5 92.3 67.6 70.5 -4.4
SFT + ER (25%) 53.2 92.1 64.5 69.9 -2.8
GRPO 57.8 92.7 72.8 74.4 -0.4
D.3 Robustness to Task Ordering
Continual learning performance is often sensitive to the task order. We evaluate RFT’s robustness by testing two distinct task orderings on both 3B and 8B models. The results in Table 9 show that the Forgetting Measure remains consistently low (ranging from -0.2% to -0.4%) regardless of the order.
Table 9: Performance evaluation under different task orderings.
Model Task Order Task 1 Task 2 Task 3 AvgAcc FM
[…]
D.4 Computational Efficiency of RIF-RFT
Regarding the computational overhead of our proposed RIF-RFT method, we provide a detailed efficiency analysis in Table 10 on $\timesH8$00 GPUs. The RIF-RFT process consists of a filtering phase (inference only) followed by training on the filtered data. Our analysis reveals that the filtering overhead is negligible (<<2% of total time) because it avoids the costly backpropagation step. Crucially, by reducing the dataset size for the subsequent training phase, RIF-RFT achieves a $\sim$44% reduction in total wall-clock time compared to standard GRPO, demonstrating that our method improves both computational and sample efficiency.
Table 10: Wall-clock time (hours) analysis comparing standard GRPO and RIF-RFT.
Dataset GRPO RIF-RFT (Train) RIF-RFT (Filter) RIF-RFT
ScienceQA 6.4 5.2 0.13 5.33
TextVQA 30.9 13.8 0.31 14.11
VizWiz 19.5 8.0 0.20 8.20
GQA 72.6 48.6 0.50 49.10
Geometry3K 2.3 0.6 0.02 0.62
PathVQA 15.4 5.6 0.15 5.75
sCLEVR 6.7 3.4 0.11 3.51
Total 153.8 85.2 1.42 86.62
D.5 Ablation on Filtering Threshold in RIF-RFT
In RIF-RFT, the filtering threshold $\tau$ determines which samples are retained for training. We use $\tau=0$ as default, meaning samples are retained if they achieve any non-zero reward across rollouts. The threshold $\tau$ controls a trade-off between data quality and quantity: higher thresholds retain only samples where the model succeeds more consistently, but this reduces the volume of training data. We conduct ablation experiments on the task sequence sCLEVR $\rightarrow$ ScienceQA $\rightarrow$ TextVQA using Qwen2.5-VL-3B. Results are presented in Table 11.
Table 11: Ablation on filtering threshold $\tau$.
$\tau$ sCLEVR SciQA TextVQA AvgAcc
0 57.2 92.5 72.1 73.9
0.1 56.8 92.1 71.4 73.4
0.2 55.9 91.6 70.2 72.6
Our default setting achieves the highest overall performance. This suggests that samples where the model has low but non-zero reward provide effective gradient signals for policy improvement. As $\tau$ increases, performance degrades across all tasks due to reduced training data volume. We recommend $\tau=0$, which maximizes the retention of informative training instances.